OpenStudy (anonymous):

Medal and Fan! Which ratio is equivalent to 6 : 7? A. 12 : 13 B. 18 : 48 C. 36 : 49 D. 48 : 56

OpenStudy (anonymous):

Simply each of them more..

OpenStudy (anonymous):

\[\frac{18}{48}\] Reduce it more..

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

Good, now reduce the following: \[\frac{36}{49}\]

OpenStudy (anonymous):

I think you can't reduce it more..

OpenStudy (anonymous):

Now we left with last one: reduce it and see what you got? \[\frac{48}{56}\]

OpenStudy (anonymous):

NOPE

OpenStudy (anonymous):

Do reduce the last one..

OpenStudy (anonymous):

The answer is D.?

OpenStudy (bohotness):

its not a because makes no sense not d but what has 36

OpenStudy (anonymous):

What will you get when you reduce 48/56??

OpenStudy (anonymous):

6/7

OpenStudy (anonymous):

6:7

OpenStudy (anonymous):

i thnk

OpenStudy (anonymous):

6/7 you got for D, so it matches with the ratio which is given in your question. So, what is the answer?